A constant 74.6-N force acts on a 4.99-kg box. The box initially is moving at 15.1 m/s in the direction of the force, and 2.93 s later the box is moving at 76.5 m/s. Determine the work (in J) done by this force. You may assume that the box experiences constant acceleration, but do not assume that only one force acts on the box.
Answer #1

The constant force of 74.6 N is the net force acting on the box and the work it does is equal to the change in the kinetic energy of the box.

Using the work-kinetic energy theorem, relate the work done by the constant force to the change in the kinetic energy of the box:

W = KEf - KEi

W = (1/2)m (Vf2 - Vi2)

W = (1/2)*4.99kg*(76.52 - 15.12 )

W = 14032.47 J